Find the area under the standard normal distribution curve. 1. Between z = 0 and z = 1.27 2. Between z = 0 and z = 0.56 3. Between z = 0 and z = -3.46 4. Between z = 0 and z = 1.54
Answer #1

Normal Probabilities (Positive) Standard Normal Probabilities Table entry Table entry for Z is the area under the standard no1/1 sic given zu N20,1) ZU N(0,1) Normal disas is synoget :: PC220)=P(Z >0)=0.5 2) Plocza 1.2%) =P (2<1.27)~P(zco) = 0.898- 0Normal Probabilities (Positive) Standard Normal Probabilities Table entry Table entry for Z is the area under the stand norma2) PCO<Z20:56) = PcZc0:56)- P(zzo) =0#123 cusing Table of software) -0.5 PCOLZ40.56) 0.2123 PCOLZCO S6)=0.2123 CS CamScannerNormal Probabilities (Positive) Standard Normal Probabilities Table entry The standard normal curve Table entry for Z is the3) RCocza PC-3.46 (220) =P (zlo) - P(Z (~3-46) 05- 0.0003 using software or -C Table) = 0.4997 P(-3.462220) P(3.46 (220) -0.410:06 A4192% Normal Probabilities (Negative) Standard Normal Probabilities Table entry Table entry for Zis the area under the4) PlocZc 1.54) = = P(ZC 1.56) - P(aco) 0.9382-05 0.4382 cusing software or Statistical Table) PC OCZ(1:54)=0.4382 PCOLZ21.54Normal Probabilities (Positive) Standard Normal Probabilities Table entry The standard normal curve Table entry for Z is the

P(Z<0)= 0